OpenStudy (anonymous):

limit (x approaches negative infinity) of ( 5^(2/x) +3^x +2)) Please help?

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

First I used t=2/x in order to get limit (t approaches 0) then i got \[Limit (t \rightarrow) 5 ^{t} +3^{\frac{ 2 }{ t }} +2 0 \]

OpenStudy (anonymous):

\[limit (t approaches 0) 5^{t} +3^{\frac{ 2 }{ t }} + 2\]

OpenStudy (anonymous):

and i know that if i seperate the limits then i could perhaps solve it. but i am stuck

OpenStudy (anonymous):

\[\lim_{t \rightarrow o} 5^t =0\] i think

OpenStudy (anonymous):

\[\lim_{t \rightarrow 0} 3^t \] i dont know

OpenStudy (zarkon):

no

OpenStudy (zarkon):

\[5^0=1\]

OpenStudy (anonymous):

\[\lim_{t \rightarrow 0} 3^\frac{ 2 }{ t}\]

OpenStudy (anonymous):

this is the part im not sure of. I mistyped earlier